The differential coefficient of f [log(x)] when f(x) log x is
A. xlog x
B. x/(log x)
C. 1/(x log x
D. (log x)/x

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
=\frac{\mathrm{d} (log(log x))}{\mathrm{d} x}
=(1/log x)(1/x)
=(1/x log x)
Hence the answer is C
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty